# EMHI: A Multimodal Egocentric Human Motion Dataset with HMD and Body-Worn IMUs

### v1

@@ Abstract
Egocentric human pose estimation (HPE) using wearable sensors is essential for VR/AR applications. Most methods rely solely on either egocentric-view images or sparse Inertial Measurement Unit (IMU) signals, leading to inaccuracies due to self-occlusion in images or the sparseness and drift of inertial sensors. Most importantly, the lack of real-world datasets containing both modalities is a major obstacle to progress in this field. To overcome the barrier, we propose EMHI, a multimodal Egocentric human Motion dataset with Head-Mounted Display (HMD) and body-worn IMUs, with all data collected under the real VR product suite. Specifically, EMHI provides synchronized stereo images from downward-sloping cameras on the headset and IMU data from body-worn sensors, along with pose annotations in SMPL format. This dataset consists of 885 sequences captured by 58 subjects performing 39 actions, totaling about 28.5 hours of recording. We evaluate the annotations by comparing them with optical marker-based SMPL fitting results. To substantiate the reliability of our dataset, we introduce MEPoser, a new baseline method for multimodal egocentric HPE, which employs a multimodal fusion encoder, temporal feature encoder, and MLP-based regression heads. The experiments on EMHI show that MEPoser outperforms existing single-modal methods and demonstrates the value of our dataset in solving the problem of egocentric HPE. We believe the release of EMHI and the method could advance the research of egocentric HPE and expedite the practical implementation of this technology in VR/AR products.
@@ 1 Introduction
Previous works [35, 47] introduced egocentric datasets using experimental fisheye camera setups to capture images and annotate 3D joints. However, these setups are impractical for real VR/AR products, which need compact, lightweight designs. Synthetic datasets [38, 2, 9] use physics engines for egocentric image rendering but suffer a domain gap with real images due to the complexity of human motion and environments. Meanwhile, the lower body may be occluded in a vision-based setting, and some body parts may fall outside the field of view (FOV) depending on the body pose. IMU-based datasets avoid occlusion but suffer from drift over time and ill-posed problems from sparse observations. Besides, existing methods [17, 52, 10] typically use synthetic IMU data from AMASS [28], which may not accurately reflect real-world noise and drift. Some datasets [39, 15] provide real 3 Degrees of Freedom (3DoF, rotation) data from XSens, while others [10] include 6DoF (rotation and position) data of head, hands, and 3DoF for lower legs, but these are small-scale and primarily used for evaluation. Recently, several large-scale multimodal datasets [27, 14] have been released, offering RGB images, upper body IMU data, and motion narrations. However, the forward-facing camera limits the egocentric view, and missing lower-body IMU signals can cause ambiguity.
Combining egocentric cameras and body-worn IMUs offers a promising multimodal solution due to their lightweight and flexible design. This configuration is also commonly found in VR scenarios. Our proposed EMHI dataset, as shown in Fig. 1, features a VR headset with two downward-sloping cameras for egocentric image capture, 6DoF head and hand tracking, and additional IMUs on an actual VR device for lower-leg 3DoF tracking. We use a markerless multi-view camera system for SMPL [25] ground truth acquirement, with accuracy and consistency refinement using IMU data, and synchronization via OptiTrack. Furthermore, we propose a new baseline method, MEPoser, integrating egocentric images and IMU data to perform real-time HPE on a standalone VR headset. The method employs a multimodal fusion encoder, a temporal feature encoder, and MLP-based regression heads to estimate SMPL body model parameters, effectively demonstrating the advantages of multimodal data fusion in enhancing pose accuracy and the value of our dataset. This approach paves the way for further research in egocentric HPE using multimodal inputs.
@@
2 Related Work
Dataset Device Real/Synth. Sensor Modality SMPL(x) Statistic
@@ 2 Related Work
DIP-IMU [Uncaptioned image] Real - Full-Body 3DoF×$\times$6 15 10 330K
FreeDancing [Uncaptioned image] Real - Full-Body 6DoF×$\times$3, 3DoF×$\times$3 - 8 532.8K
Nymeria [Uncaptioned image] Real Binocular Forward-Facing Upper-Body 6DoF×$\times$3 20 264 260M
Ego-Exo4D (Ego Pose) [Uncaptioned image] Real Binocular Forward-Facing Head 6DoF×$\times$1 - - - 9.6M
Ours [Uncaptioned image] Real Binocular Downward-Sloping Full-Body 6DoF×$\times$3, 3DoF×$\times$2 39 58 3.07M
@@
2.1 Egocentric Motion Dataset
@@
Unlike previous HPE datasets captured by third-view cameras [16, 3, 19, 50, 5], vision-based egocentric motion datasets provide first-person perspective images using head-mounted monocular or binocular cameras, with the corresponding annotations of the wearer’s poses. Mo2Cap2 [47] and xR-EgoPose [38] made an early effort to build the synthetic monocular dataset with a downward-facing fisheye camera. EgoPW [43] proposed the first in-the-wild dataset and was followed by EgoGTA [44] and ECHA [23] with the same camera setting. EgoWholeBody [45] is the latest synthetic dataset providing high-quality images and SMPL-X annotations. EgoCap [35] is a pioneer binocular dataset captured by helmet-mounted stereo cameras, containing 30K frames recorded in a lab environment. EgoGlass [51] optimized the binocular setup with two front-facing cameras mounted on the glasses frames. To relieve the dataset scale limitation, UnrealEgo [1] proposed a large-scale and highly realistic stereo synthetic dataset with 450K stereo views and was extended to 1.25M in UnrealEgo2 [2]. SynthEgo [9] extended synthetic datasets with more identities and environments, annotated with SMPL-H for better body shape descriptions.
Sparse IMU-based datasets provide an alternative for this problem. AMASS [28] can provide a large-scale synthetic IMU dataset. TotalCapture [39] and DIP-IMU [15] offered real IMU data captured by Xsens and SMPL pose annotations obtained by marker-based optical mocap system and IMU-based method [41] respectively. PICO-FreeDancing [10] provided sparse IMU data with SMPL format GT fitting using OptiTrack data. However, real IMU-based datasets are generally limited in scale.
Multimodal datasets[11, 13, 8, 34] have attracted significant attention in recent years due to the complementarity of different data modalities. Ego-Exo4D [14], Nymeria [27] and SimXR [26] captured real-world images by Project Aria glasses [36], along with the IMU data in upper-body. Ego-Exo4D provided up to 9.6M image frames with annotations of the body and hand joint positions. Nymeria further offered SMPL format data derived from Xsens mocap suits, with limited clothing diversity of the captured body. However, in these datasets, either the forward-facing perspective restricts the perception range of the wearer’s body, or they have not integrated downward-sloping perspectives and sparse full-body IMU signals on actual VR/AR devices.
@@
2.2 Egocentric Human Pose Estimation Methods
Existing egocentric HPE methods primarily utilize single-modal input. Vision-based methods have been widely investigated  [35, 43, 44, 22, 23, 20, 30, 29, 9]. Jian Wang et al. [42] introduced a spatiotemporal optimization method for single-view egocentric sequences to obtain the 3D skeleton results. Recently, Wang et al. [45] proposed an egocentric motion capture method that combines the vision transformer for undistorted image patch feature extracting and uses diffusion-based motion priors for pose refinement. However, 3D pose estimation from a single image remains challenging due to the lack of depth information. To address this, UnrealEgo [1] introduced stereo egocentric skeleton tracking methods by integrating a weight-shared encoder for stereo heatmap generation and a multi-branch autoencoder for 3D pose prediction. Hiroyasu Akada et al. [2] further enhanced this with a transformer-based model utilizing 3D scene information and temporal features. Despite these advancements, challenges with invisible body parts due to self-occlusion and out-of-view joints persist.
Methods using sparse tracking signals from body-worn IMUs have garnered significant attentions [46, 40, 15, 48, 49, 18, 12, 7]. In egocentric VR and AR scenarios, there are inherently three 6DoF tracking points for the head and hands, with the option to add two additional 3DoF IMUs on the legs. AvatarPoser [17] proposed a global pose prediction framework combining transformer structures with inverse kinematics (IK) optimization, while AvatarJLM [52] introduced a two-stage approach that models joint-level features and uses them as spatiotemporal transformer tokens to achieve smooth action capture. HMD-Poser [10] integrated these inputs, presenting a lightweight temporal-spatial learning method for full-body global 6DoF body action recovery. However, IMU-based data faces challenges such as drift and sparsity. Many current methods rely on synthetic IMU data from AMASS [28], which often fails to capture real-world noise and drift accurately, leading to overfitting as models are not exposed to the complexities and imperfections of real-world conditions.
3 EMHI Dataset
@@
3.1 Data Capture System
3.1.1 Hardware
@@
3.1.2 Temporal Synchronization
Kinect and Optitrack systems rely on a signal transmitter device to trigger simultaneously, ensuring their inter-frame alignment. EgoSensorKit and Optitrack could also be synchronized offline with the headset IMU’s and ORB’s angular velocity, according to the motion correlation method in [32]. Finally, the data frames of Kinect and EgoSensorKit are aligned via Optitrack as a bridge. As the recorded frame rate is 30Hz, the maximum synchronization deviation will reach up to 16.5 ms which might be notable during fast motion. So the annotations are further post-processed with linear interpolation to better align with the EgoSensorKit’s timestamps.
3.1.3 Spatial Alignment
@@
3.2 Ground-Truth Acquisition
3.2.1 Keypoints Annotation
We use HRNet [37] to detect 2D keypoints in multi-view Kinect RGB images with the body25 format [6]. Then, we follow HuMMan [5] to derive 3D keypoints annotations $P_{3D}$ by triangulation with camera parameters obtained in the spatial alignment, in which we also import smoothness and bone length constraints for $P_{3D}$ to reduce temporal jitter and improve human shape consistency.
3.2.2 SMPL Fitting
@@
where $\theta\in\mathbb{R}^{75}$ and $\beta\in\mathbb{R}^{10}$ are optimized SMPL pose and shape parameters and $\lambda_{*}$ are balance weight (see Appendix). For occluded joints, we introduce the rotation term to encourage the pose consistency with transferred IMU data as follows:
$E_{rot}=\sum_{j}||\mathcal{F}(\theta)_{j}-R_{j}||,$ (2)
where $j\in\{head,wrist,knee\}$ and $\mathcal{F}$ indicate a forward kinematic (FK) to get the joint global rotation. Other energy terms are like previous works [4, 5, 31], in which $E_{joint}$ minimizes the 3D distance between $P_{3D}$ and regressed SMPL joints. $E_{prior}(\theta)$ is Vposer prior from SMPLify-X. $E_{smooth}$ helps to keep smooth pose tracking, while the shape regularization term $E_{reg}$ penalizes large shape variance. With the space alignment result, the SMPL results could transfer from world space to the egocentric camera coordinate and obtain 2D pose annotations on egocentric images.
4 A New Baseline Method: MEPoser
@@
4.1 Multimodal Fusion Encoder
The multimodal encoder first has separate feature encoders for different modalities, i.e., two weight-sharing CNN backbones for images and an MLP network for IMU data. To make MEPoser run in real-time on HMD, we use a lightweight RegNetY-400MF [33] backbone, which takes stereo images $\{Img_{left}^{t_{i}},Img_{right}^{t_{i}}\}\in\mathbb{R}^{640\times 480\times 1}$ as inputs, and generates 2D image features represented as $\{\mathrm{F}_{left}^{t_{i}},\mathrm{F}_{right}^{t_{i}}\}\in\mathbb{R}^{80\times 60\times 256}$. These features are then concatenated and forwarded to a few convolution layers to infer a set of heatmaps $\{\mathrm{H}_{left}^{t_{i}},\mathrm{H}_{right}^{t_{i}}\}\in\mathbb{R}^{80\times 60\times J}$. Here we predict 22 joints of the SMPL, i.e., $J=22$. To train the RegNetY-400MF backbone, we calculate the binary cross-entropy with logits loss (BCEWithLogitsLoss) $\mathcal{L}_{heatmap}^{2D}$ between the GT heatmaps and the estimated 2D heatmaps. Then, the predicted heatmaps are flattened and forwarded to an MLP network to obtain the image feature. We have obtained the IMU and image features so far. To boost the performance of the pose estimation, we added a 3D module to estimate the 3D joint positions in both the local camera coordinate and the global world coordinate. Specifically, given the image features from stereo heatmaps, an MLP network first encodes them to estimate the 3D joint positions in the local camera coordinate $\mathrm{\hat{P}}_{local}^{J\times 3}$. Then, these joints are transferred to the global SMPL coordinate $\mathrm{\hat{P}}_{global}^{J\times 3}$ with the offline calibration results and the online headset’s 6DoF data. $\mathrm{\hat{P}}_{local}^{J\times 3}$ and $\mathrm{\hat{P}}_{global}^{J\times 3}$ are used to calculate the 3D joint loss $\mathcal{L}_{joints}^{Local}$ and $\mathcal{L}_{joints}^{Global}$, respectively. Next, the joint positions $\mathrm{\hat{P}}_{global}^{J\times 3}$ are flattened and forwarded to an MLP network to obtain the 3D joint features. Finally, the IMU, image, and 3D joint features are concatenated to output the multimodal fused feature $f^{t_{i}}$.
4.1.1 Temporal Feature Encoder
As demonstrated in HMD-Poser [10], temporal correlation information is the key to tracking accurate human motions. However, the multimodal fused features $\{f^{t_{i}}\}$ are still temporally isolated. To solve this problem, Transformer and RNN are adopted in existing methods. Although Transformer-based methods [52] have achieved state-of-the-art results in HPE, their computational costs are much higher than those of RNN-based methods. To ensure our method runs in real-time on HMDs, we introduce a lightweight LSTM-based temporal feature encoder. Specifically, the encoder is composed of a stack of $N=3$ identical blocks. And each block has two sub-layers. The first is an LSTM module to learn the temporal representation, and the second is a simple fully connected feed-forward network. We employ a residual connection followed by layer normalization.
4.2 SMPL Decoder
4.3 Training MEPoser
For the overall training loss, we combine a smooth loss $\mathcal{L}_{smooth}$ with the above losses, including 2D heatmap loss $\mathcal{L}_{heatmap}^{2D}$, 3D joint loss $\mathcal{L}_{joints}^{Local}$, $\mathcal{L}_{joints}^{Global}$ and SMPL loss $\mathcal{L}_{SMPL}$. The smooth loss from HMD-Poser [10] is adopted to further enhance the temporal smoothness.
where $\lambda_{*}$ here are the balance weights (see Appendix).
@@ 4.3 Training MEPoser
MEPoser(Ours) 4.1 3.7 2.5 2.7 5.1 3.2 161.8
@@
MEPoser(Ours) 4.7 4.8 2.9 3.2 7.0 3.8 204.9
@@
5 Experiment
5.1 Dataset Splitting
@@
5.2 Comparison
Our multimodal dataset fills a gap in egocentric VR scenarios. To validate the dataset and the corresponding baseline methods, we conducted the following comparison using our dataset between MEPoser against the latest single-modal methods: UnrealEgo [1] with stereo egocentric images as inputs and HMD-Poser [10] with IMU observations from HMD and 2 leg motion trackers.
The quantitative results in Tab. 2 show that MEPoser outperforms existing single-modal methods for egocentric HPE. Compared to UnrealEgo, our method reduces the MPJPE (Mean Per Joint Position Error, cm) by 32.7% and 25% in protocol1 and protocol2, respectively. Notably, MEPoser significantly enhances the smoothness of estimation results by using a temporal LSTM structure. In comparison with HMD-Poser, MEPoser shows 36.2% and 31.4% reduction in MPJPE on two test sets by combining the egocentric image features and slight improvement of joint rotation accuracy according to the MPJRE (Mean Per-Joint Rotation Error, ) results. The results also demonstrate the improved generalizability of MEPoser across subjects and actions, validating the effectiveness of the multimodal setting and the value of our dataset in solving egocentric HPE.
@@ 5.2 Comparison
5.3 Annotation Cross-validation
To further assess the accuracy of our dataset annotations, we randomly capture 9 motion sequences for cross-validation by simultaneously using both our system and an optical marker-based motion capture system. The subject wears a suit attached with reflective markers tracked by the OptiTrack system. Marker-based SMPL parameters are then derived from the MoCap data using MoSh++ [24], with a temporal filter applied to reduce jittering. As shown in Tab. 3, the low error metrics and variance demonstrate the robustness of our annotation pipeline and the high quality of the dataset. Besides the validation, every sequence of our dataset has been inspected manually to eliminate the data with erroneous annotations.
@@ 5.3 Annotation Cross-validation
6 Conclusion

### v2

@@ Abstract
Egocentric human pose estimation (HPE) using wearable sensors is essential for VR/AR applications. Most methods rely solely on either egocentric-view images or sparse Inertial Measurement Unit (IMU) signals, leading to inaccuracies due to self-occlusion in images or the sparseness and drift of inertial sensors. Most importantly, the lack of real-world datasets containing both modalities is a major obstacle to progress in this field. To overcome the barrier, we propose EMHI, a multimodal Egocentric human Motion dataset with Head-Mounted Display (HMD) and body-worn IMUs, with all data collected under the real VR product suite. Specifically, EMHI provides synchronized stereo images from downward-sloping cameras on the headset and IMU data from body-worn sensors, along with pose annotations in SMPL format. This dataset consists of 885 sequences captured by 58 subjects performing 39 actions, totaling about 28.5 hours of recording. We evaluate the annotations by comparing them with optical marker-based SMPL fitting results. To substantiate the reliability of our dataset, we introduce MEPoser, a new baseline method for multimodal egocentric HPE, which employs a multimodal fusion encoder, temporal feature encoder, and MLP-based regression heads. The experiments on EMHI show that MEPoser outperforms existing single-modal methods and demonstrates the value of our dataset in solving the problem of egocentric HPE. We believe the release of EMHI and the method could advance the research of egocentric HPE and expedite the practical implementation of this technology in VR/AR products. Project page at: https://pico-ai-team.github.io/EMHI/
Introduction
@@
Previous works (rhodin2016egocap; xu2019mo) introduced egocentric datasets using experimental fisheye camera setups to capture images and annotate 3D joints. However, these setups are impractical for real VR/AR products, which need compact, lightweight designs. Synthetic datasets (tome2019xr; akada20243d; cuevas2024simpleego) use physics engines for egocentric image rendering but suffer a domain gap with real images due to the complexity of human motion and environments. Meanwhile, the lower body may be occluded in a vision-based setting, and some body parts may fall outside the field of view (FOV) depending on the body pose. IMU-based datasets avoid occlusion but suffer from drift over time and ill-posed problems from sparse observations. Besides, existing methods (jiang2022avatarposer; zheng2023realistic; dai2024hmd) typically use synthetic IMU data from AMASS (mahmood2019amass), which may not accurately reflect real-world noise and drift. Some datasets (trumble2017total; huang2018deep) provide real 3 Degrees of Freedom (3DoF, rotation) data from XSens, while others (dai2024hmd) include 6DoF (rotation and position) data of head, hands, and 3DoF for lower legs, but these are small-scale and primarily used for evaluation. Recently, several large-scale multimodal datasets (ma2024nymeria; grauman2024ego) have been released, offering RGB images, upper body IMU data, and motion narrations. However, the forward-facing camera limits the egocentric view, and missing lower-body IMU signals can cause ambiguity.
Combining egocentric cameras and body-worn IMUs offers a promising multimodal solution due to their lightweight and flexible design. This configuration is also commonly found in VR scenarios. Our proposed EMHI dataset, as shown in Fig. 1, features a VR headset with two downward-sloping cameras for egocentric image capture, 6DoF head and hand tracking, and additional IMUs on an actual VR device for lower-leg 3DoF tracking. We use a markerless multi-view camera system for SMPL (SMPL:2015) ground truth acquirement, with accuracy and consistency refinement using IMU data, and synchronization via OptiTrack. Furthermore, we propose a new baseline method, MEPoser, integrating egocentric images and IMU data to perform real-time HPE on a standalone VR headset. The method employs a multimodal fusion encoder, a temporal feature encoder, and MLP-based regression heads to estimate SMPL body model parameters, effectively demonstrating the advantages of multimodal data fusion in enhancing pose accuracy and the value of our dataset. This approach paves the way for further research in egocentric HPE using multimodal inputs.
@@
Related Work
Dataset Device Real/Synth Sensor Modality SMPL(x) Statistic
@@
[…]
@@
Egocentric Motion Dataset
@@
Vision-based egocentric motion datasets (zhao2021egoglass; wang2022estimating; wang2023scene; liuegofish3d) provide first-person perspective images using head-mounted cameras, with the corresponding annotations of the wearer’s poses. Mo2Cap2 (xu2019mo) and xR-EgoPose (tome2019xr) made an early effort to build the synthetic monocular dataset with a downward-facing fisheye camera. EgoWholeBody (wang2024egocentric) is the latest synthetic dataset providing high-quality images and SMPL-X annotations. EgoCap (rhodin2016egocap) is a pioneer binocular dataset captured by helmet-mounted stereo cameras, containing 30K frames recorded in a lab environment. To relieve the dataset scale limitation, UnrealEgo (akada2022unrealego) proposed a large-scale and highly realistic stereo synthetic dataset with 450K stereo views and was extended to 1.25M in UnrealEgo2 (akada20243d). SynthEgo (cuevas2024simpleego) extended synthetic datasets with more identities and environments, annotated with SMPL-H for better body shape descriptions.
Sparse IMU-based datasets provide an alternative for this problem. AMASS (mahmood2019amass) can provide a large-scale synthetic IMU dataset. TotalCapture (trumble2017total) and DIP-IMU (huang2018deep) offered real IMU data captured by Xsens and SMPL pose annotations obtained by marker-based optical mocap system and IMU-based method (von2017sparse) respectively. PICO-FreeDancing (dai2024hmd) provided sparse IMU data with SMPL format GT fitting using OptiTrack data.
Multimodal datasets(damen2022rescaling; gong2023mmg; cha2021mobile; rai2021home) have attracted significant attention in recent years due to the complementarity of different data modalities. Ego-Exo4D (grauman2024ego), Nymeria (ma2024nymeria) and SimXR (luo2024real) captured real-world images by Project Aria glasses (somasundaram2023project), along with the IMU data in upper-body. Ego-Exo4D provided up to 9.6M image frames with annotations of the body and hand joint positions. Nymeria further offered SMPL format data derived from Xsens mocap suits, with limited clothing diversity of the captured body. However, in these datasets, either the forward-facing perspective restricts the perception range of the wearer’s body, or they have not integrated downward-sloping perspectives and sparse full-body IMU signals on actual VR/AR devices.
@@
Egocentric Human Pose Estimation Methods
Vision-based methods have been widely investigated  (rhodin2016egocap; Wang_2021_ICCV; wang2022estimating; wang2023scene; akada2022unrealego; liu2022ego+; liuegofish3d; kang2023ego3dpose; park2023domain; cuevas2024simpleego). For single egocentric pose estimation, Wang et al. (wang2024egocentric) proposed an egocentric motion capture method that combines the vision transformer for undistorted image patch feature extracting and uses diffusion-based motion priors for pose refinement. However, 3D pose estimation from a single image remains challenging due to the lack of depth information. To address this, EgoPoseFormer (yang2024egoposeformer) introduced stereo egocentric skeleton tracking methods by a two-stage transformer-based corse-to-fine pose estimator for 3D pose prediction. Hiroyasu Akada et al. (akada20243d) further enhanced this with a transformer-based model utilizing 3D scene information and temporal features. Despite these advancements, challenges with invisible body parts due to self-occlusion and out-of-view joints persist.
Methods using sparse tracking signals from body-worn IMUs have garnered significant attentions (winkler2022questsim; SIP; yi2021transpose; yi2022physical; TIP; AGROL).In egocentric VR and AR scenarios, there are inherently three 6DoF tracking points for the head and hands, with the option to add two additional 3DoF IMUs on the legs. AvatarPoser (jiang2022avatarposer) proposed a global pose prediction framework combining transformer structures with inverse kinematics (IK) optimization, while AvatarJLM (zheng2023realistic) introduced a two-stage approach that models joint-level features and uses them as spatiotemporal transformer tokens to achieve smooth action capture. HMD-Poser (dai2024hmd) integrated these inputs, presenting a lightweight temporal-spatial learning method for full-body global 6DoF body action recovery. However, IMU-based data faces challenges such as drift and sparsity.
EMHI Dataset
@@
Data Capture System
Hardware
@@
Temporal Synchronization
Kinect and Optitrack systems rely on a signal transmitter device to trigger simultaneously, ensuring their inter-frame alignment. EgoSensorKit and Optitrack could also be synchronized offline with the headset IMU’s and ORB’s angular velocity, according to the motion correlation method in (qiu2020real). Finally, the data frames of Kinect and EgoSensorKit are aligned via Optitrack as a bridge. As the recorded frame rate is 30Hz, the maximum synchronization deviation will reach up to 16.5 ms which might be notable during fast motion. So the annotations are further post-processed with linear interpolation to better align with the EgoSensorKit’s timestamps.
Spatial Alignment
@@
Ground-Truth Acquisition
Keypoints Annotation
We use HRNet (sun2019deep) to detect 2D keypoints in multi-view Kinect RGB images with the body25 format (cao2017realtime). Then, we follow HuMMan (cai2022humman) to derive 3D keypoints annotations $P_{3D}$ by triangulation with camera parameters obtained in the spatial alignment, in which we also import smoothness and bone length constraints for $P_{3D}$ to reduce temporal jitter and improve human shape consistency.
SMPL Fitting
@@
where $\theta\in\mathbb{R}^{75}$ and $\beta\in\mathbb{R}^{10}$ are optimized SMPL pose and shape parameters and $\lambda_{*}$ are balance weight. For occluded joints, we introduce the rotation term to encourage the pose consistency with transferred IMU data as follows:
[…]
where $j\in\{head,wrist,knee\}$ and $\mathcal{F}$ indicate a forward kinematic (FK) to get the joint global rotation. Other energy terms are like previous works (bogo2016keep; cai2022humman; pavlakos2019expressive), in which $E_{joint}$ minimizes the 3D distance between $P_{3D}$ and regressed SMPL joints. $E_{prior}(\theta)$ is Vposer prior from SMPLify-X. $E_{smooth}$ helps to keep smooth pose tracking, while the shape regularization term $E_{reg}$ penalizes large shape variance. With the space alignment result, the SMPL results could transfer from world space to the egocentric camera coordinate and obtain 2D pose annotations on egocentric images.
A New Baseline Method: MEPoser
@@
MEPoser-CV 5.4 4.5 2.9 3.3 6.3 3.8 511.0
MEPoser-IMU 5.0 6.2 3.6 5.0 8.0 5.2 121.7
MEPoser-Full 4.1 3.7 2.5 2.7 5.1 3.2 161.8
@@
MEPoser-CV 6.0 5.4 3.5 3.8 7.8 4.4 566.5
MEPoser-IMU 5.7 7.1 4.2 5.4 9.9 6.6 161.7
MEPoser-Full 4.7 4.8 2.9 3.2 7.0 3.8 204.9
@@
Multimodal Fusion Encoder
The multimodal encoder first has separate feature encoders for different modalities, i.e., two weight-sharing CNN backbones for images and an MLP network for IMU data. To make MEPoser run in real-time on HMD, we use a lightweight RegNetY-400MF (radosavovic2020designing) backbone, which takes stereo images $\{Img_{left}^{t_{i}},Img_{right}^{t_{i}}\}\in\mathbb{R}^{640\times 480\times 1}$ as inputs, and generates 2D image features represented as $\{\mathrm{F}_{left}^{t_{i}},\mathrm{F}_{right}^{t_{i}}\}\in\mathbb{R}^{80\times 60\times 256}$. These features are then concatenated and forwarded to a few convolution layers to infer a set of heatmaps $\{\mathrm{H}_{left}^{t_{i}},\mathrm{H}_{right}^{t_{i}}\}\in\mathbb{R}^{80\times 60\times J}$. Here we predict 22 joints of the SMPL, i.e., $J=22$. To train the RegNetY-400MF backbone, we calculate the binary cross-entropy with logits loss (BCEWithLogitsLoss) $\mathcal{L}_{heatmap}^{2D}$ between the GT heatmaps and the estimated 2D heatmaps. Then, the predicted heatmaps are flattened and forwarded to an MLP network to obtain the image feature. We have obtained the IMU and image features so far. To boost the performance of the pose estimation, we added a 3D module to estimate the 3D joint positions in both the local camera coordinate and the global world coordinate. Specifically, given the image features from stereo heatmaps, an MLP network first encodes them to estimate the 3D joint positions in the local camera coordinate $\mathrm{\hat{P}}_{local}^{J\times 3}$. Then, these joints are transferred to the global SMPL coordinate $\mathrm{\hat{P}}_{global}^{J\times 3}$ with the offline calibration results and the online headset’s 6DoF data. $\mathrm{\hat{P}}_{local}^{J\times 3}$ and $\mathrm{\hat{P}}_{global}^{J\times 3}$ are used to calculate the 3D joint loss $\mathcal{L}_{joints}^{Local}$ and $\mathcal{L}_{joints}^{Global}$, respectively. Next, the joint positions $\mathrm{\hat{P}}_{global}^{J\times 3}$ are flattened and forwarded to an MLP network to obtain the 3D joint features. Finally, the IMU, image, and 3D joint features are concatenated to output the multimodal fused feature $f^{t_{i}}$.
Temporal Feature Encoder
As demonstrated in HMD-Poser (dai2024hmd), temporal correlation information is the key to tracking accurate human motions. However, the multimodal fused features $\{f^{t_{i}}\}$ are still temporally isolated. To solve this problem, Transformer and RNN are adopted in existing methods. Although Transformer-based methods (zheng2023realistic) have achieved state-of-the-art results in HPE, their computational costs are much higher than those of RNN-based methods. To ensure our method runs in real-time on HMDs, we introduce a lightweight LSTM-based temporal feature encoder. Specifically, the encoder is composed of a stack of $N=3$ identical blocks. And each block has two sub-layers. The first is an LSTM module to learn the temporal representation, and the second is a simple fully connected feed-forward network. We employ a residual connection followed by layer normalization.
SMPL Decoder
Training MEPoser
For the overall training loss, we combine a smooth loss $\mathcal{L}_{smooth}$ with the above losses, including 2D heatmap loss $\mathcal{L}_{heatmap}^{2D}$, 3D joint loss $\mathcal{L}_{joints}^{Local}$, $\mathcal{L}_{joints}^{Global}$ and SMPL loss $\mathcal{L}_{SMPL}$. The smooth loss from HMD-Poser (dai2024hmd) is adopted to further enhance the temporal smoothness.
Experiment
Dataset Splitting
@@
Comparison
To validate the dataset and the corresponding baseline methods, we conducted comparisons using our dataset between MEPoser against the latest single-modal methods. The quantitative results in Tab. 2 show that MEPoser outperforms existing single-modal methods for egocentric HPE. Compared to Unrealego (akada2022unrealego) which takes stereo egocentric images as inputs, our method reduces the MPJPE (Mean Per Joint Position Error, cm) by 32.7% and 25% in protocol1 and protocol2, respectively. Notably, MEPoser significantly enhances the smoothness of estimation results by using a temporal LSTM structure. Besides, our method obtains 27.5% and 23.8% enhancement in joint location precious in contrast to EgoPoseFormer (yang2024egoposeformer). In comparison with whole-body IMU-based method HMD-Poser (dai2024hmd), MEPoser shows 36.2% and 31.4% reduction in MPJPE on two test sets by combining the egocentric image features and slight improvement of joint rotation accuracy according to the MPJRE (Mean Per-Joint Rotation Error, ). The results also demonstrate the improved generalizability of MEPoser across subjects and actions, validating the effectiveness of the multimodal setting and the value of our dataset in solving egocentric HPE. We also include ablation experiments to investigate the impact of different components in our network and the results demonstrate that each modality contributes to performance improvements.
@@
Annotation Cross-validation
@@
To further assess the accuracy of our dataset annotations, we randomly capture 9 motion sequences for cross-validation by simultaneously using both our system and an optical marker-based motion capture system. The subject wears a suit attached with reflective markers tracked by the OptiTrack system. Marker-based SMPL parameters are then derived from the MoCap data using MoSh++ (loper2014mosh), with a temporal filter applied to reduce jittering. As shown in Tab. 3, the low error metrics and variance demonstrate the robustness of our annotation pipeline and the high quality of the dataset. Besides the validation, every sequence of our dataset has been inspected manually to eliminate the data with erroneous annotations.
Conclusion